OpenStudy (theedwardsfamily):

Numbers p and q are negative integers. Which statements are always true? Choose all answers that are correct.

OpenStudy (michele_laino):

where are your option, please?

OpenStudy (michele_laino):

oops..options...

OpenStudy (theedwardsfamily):

sorry

OpenStudy (theedwardsfamily):

@Michele_Laino

OpenStudy (michele_laino):

I think A is correct, since if p=-9 and q=-5, then p+q=-14 and, in general C is false, for example: if p=-1 and q=-29, then p-q=-1-(-29)=28

OpenStudy (michele_laino):

whereas B is always true

OpenStudy (michele_laino):

and D is always false

OpenStudy (michele_laino):

reassuming D is false, since it contradicts the sign rule of algebra

OpenStudy (theedwardsfamily):

ok thanks again @Michele_Laino !

OpenStudy (michele_laino):

thanks! @TheEdwardsFamily